OpenStudy (anonymous):

if f(x) = −2x − 4 and g(x) = x + 4 find (f/g)(-8)

OpenStudy (anonymous):

If I am reading the problem correctly, it is asking for f(-8)/g(-8). if so, plug in -8 to each function and then divide

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

To do this, simply find \[\frac{f(x)}{g(x)}\] and plug in the value. It should be -3

OpenStudy (anonymous):

(-2x-4)/(x+4) substitute x=8 (-16-4)/8+4) =-20/12

OpenStudy (anonymous):

substitute x = -8